OpenStudy (anonymous):

how many minutes would 1/12 of a half day be

OpenStudy (anonymous):

well 1/12 is equal to 2/24 and there are 24 hours in a day so 2/24 would be equal to 2 hours therefore 120 minutes

jimthompson5910 (jim_thompson5910):

there are 24 hrs in a day, so 1/12 of that is (1/12)*24 = 24/12 = 2 hours and 2 hrs = 2*60 = 120 minutes So 1/12 of a day is 120 minutes